# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

### v1

@@ 3.2 A Unified Information Theoretic Framework
Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a minimal sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
In pure statistical terms, Definition 3.1 implies that an ideal representation $\bm{Z}$ is the simplest mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$. We therefore assume the following dependency structures in terms of directed graphical models:
Assumption 3.2.
The dependency graphs of COMRL are given by
satisfying $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$. To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is fully captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$. We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$. Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be maximally causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component. Accordingly, we name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively. With this prior knowledge and assumption, we present the central theorem of this paper:
@@
Let $\equiv$ denote equality up to a constant,
$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\text{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\underbrace{I(\bm{Z};\bm{X})}_{\text{primary + lesser causality}}$
@@
  2. 2.
@@
  3. 3.
@@
See Appendix B, Theorem B.1. ∎
Theorem 3.3 induces several key observations. Firstly, the FOCAL and CORRO objective operates as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively. Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 3.1, CORRO, which maximizes a lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical guarantees. However, FOCAL which maximizes an upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ is not guaranteed to optimize $I(\bm{Z};\bm{M})$. Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may also increase the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is not ideal since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$. This explains why FOCAL is less robust to context shift compared to CORRO.
Secondly, CSRO as a latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives. In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which may form a tighter bound on the real objective $I(\bm{Z};\bm{M})$. This explains the improved performance of CSRO compared to FOCAL and CORRO.
3.3 UNICORN: Implementation of The Information Bottleneck Principle on COMRL
Given Definition 3.1 of the task representation learning, we now further generalize the pre-existing COMRL algorithms based on the information bottleneck principle (Tishby et al., 2000; Tishby and Zaslavsky, 2015). Namely, finding an optimal task representation $\bm{Z}$ can be formulated as the minimization of the following Lagrangian
$\mathcal{L}_{IB}=I(\bm{Z};\bm{X})-\beta I(\bm{Z};\bm{M})$ (7)
subject to the Markov chain constraint. The positive Lagrange multiplier $\beta$ acts as a tradeoff parameter between the complexity/rate of the representation, $R=I(\bm{Z};\bm{X})$, and the amount of preserved relevant information, $I_{M}=I(\bm{Z};\bm{M})$. For implementation, the first term is just FOCAL loss by Theorem 3.3, whereas we propose to approximate the second term as a convex combination of the FOCAL and CORRO loss, analogous to CSRO:
@@ 3.3 UNICORN: Implementation of The Information Bottleneck Principle on COMRL
Figure 2: Meta-training procedure of UNICORN. Built upon the classic COMRL methods FOCAL, UNICORN implements the information bottleneck principle by jointly optimizing the FOCAL loss and a reconstruction loss, using a variational auto-encoder to obtain robust task representation $\bm{z}$.
where $0\leq\alpha\leq 1$. Substituting Eqn 8 into 7, we may rescale $\mathcal{L}_{IB}$ as
$\mathcal{L}_{IB}=-\left(I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (9)
For implementation of $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, however, it would require additional approximations such as generating negative samples for contrastive learning using a trained model in CORRO, or using a CLUB upper bound of the mutual information in CSRO. Instead, we propose to rewrite it according to the chain rule of mutual information (Yeung, 2008):
@@
$I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ $=\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log\frac{p(\bm{x_{t}}|\bm{z},\bm{x_{b}})}{p(\bm{x_{t}})}$
$\equiv\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log p(\bm{x_{t}}|\bm{z},\bm{x_{b}})$
$\simeq\int_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\underbrace{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}_{\text{encoder}}\underbrace{\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})}_{\text{decoder}}$
$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}\left[\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right],$ (11)
which induces an unbiased estimator of $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$, reminiscent of VAE (Kingma and Welling, 2014). We therefore implement the objective of our proposed algorithm UNICORN as a linear combination of this reconstruction loss $\mathcal{L}_{\text{recon}}:=-I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ and the FOCAL loss $\mathcal{L}_{\text{FOCAL}}=-I(\bm{Z};\bm{X})$:
@@
$\mathcal{L}_{\text{UNICORN}}$ $:=\mathcal{L}_{IB}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
$=-\left(I(\bm{X}_{t};\bm{Z},\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (12)
UNICORN can be interpreted as a generalization of the FOCAL and CORRO objecitves by effectively striking a balance between the primary causality $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and the lesser causality $I(\bm{Z};\bm{X}_{b})$. Figure 2 gives an overview of our framework, which enjoys three core advantages:
  • Robustness and bias-reduction. By optimizing a reconstruction loss $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ equivalent to the CORRO objective $I(\bm{Z};\bm{X_{t}}|\bm{X_{b}})$, UNICORN mitigates the spurious correlation in FOCAL, leading to more robust task representation $\bm{Z}$. It is also verified in Section 4.5 that UNICORN is robust to a wide range of data qualities. Moreover, the unbiased estimator in Eqn 11 circumvents additional approximations as in CORRO and CSRO, making UNICORN a bias-reduced and improved version of previous SOTAs.
  • Broader applicability domain. Unlike FOCAL/CORRO which assumes uniform task distribution, the unbiased estimator in Eqn 11 can be efficiently implemented by Monte Carlo sampling on offline datasets drawn from any distribution. Additionally, we will show in Section 5.1 that UNICORN is transferable to other model architectures such as Decision Transformer (DT).
  • General and unified paradigm. Derived from the general information bottleneck principle, our framework encompasses a series of milestones in COMRL including FOCAL, CORRO, CSRO and a concurrent method GENTLE (Zhou et al., 2023), which solely optimizes the reconstruction loss with a similar implementation. A unified picture of these pre-existing methods being special cases of UNICORN in the $\alpha$-$\beta$ plane is depicted in Figure 3. A hyper-parameter study showing the effect of the UNICORN feasible region is presented in Appendix C.1. Moreover, the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ can operate as a model of the MDP to generate additional data using customized latent $\bm{z}$, which is explored in Section 5.2. In summary, UNICORN not only provides a unified framework of model-free COMRL, it may also be exploited and lead to novel model-based algorithms.

### v2

@@ 2.2 A Unified Information Theoretic Framework
Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
In pure statistical terms, Definition 2.1 implies that an ideal representation $\bm{Z}$ is a mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$. We therefore define the following dependency structures in terms of directed graphical models:
Definition 2.2.
The dependency graphs of COMRL is given by Fig. 2, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
Figure 2: Graphical Models of COMRL.
For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$. To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is primarily captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$. The only exception is when tasks differ in initial state distribution $\rho_{0}$ or transition dynamics $T$, in which case the state variable $\bm{S}$ also depends on $\bm{M}$. We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$. Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be maximally causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component. Accordingly, we name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively. With this prior knowledge and assumption, we present the central theorem of this paper:
@@
Let $\equiv$ denote equality up to a constant, then
$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\textup{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\quad I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})=\underbrace{I(\bm{Z};\bm{X})}_{\textup{primary + lesser causality}}$
@@
    $\mathcal{L}_{\textup{FOCAL}}\equiv-I(\bm{Z};\bm{X})$.
  2. 2.
    $\mathcal{L}_{\textup{CORRO}}\equiv-I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
  3. 3.
    $\mathcal{L}_{\textup{CSRO}}\geq-\left((1-\lambda)I(\bm{Z};\bm{X})+\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\right)$.
Proof.
See Appendix B. ∎
Theorem 2.3 induces several key observations. Firstly, the FOCAL and CORRO objectives operate as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively. Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 2.1, CORRO, which maximizes the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical assurance. However, FOCAL which maximizes the upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ provides no guarantee for $I(\bm{Z};\bm{M})$. Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may instead significantly elevates the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$. This explains why FOCAL is less robust to context shift compared to CORRO.
Secondly, CSRO as the latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives. In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T\,\&\,\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$. This accounts for the improved performance of CSRO compared to FOCAL and CORRO.
2.3 Instantiations of UNICORN
By providing a unified view of pre-existing COMRL algorithms, Theorem 2.3 opens up avenues for novel algorithmic implementations by seeking alternative approximations of the true objective $I(\bm{Z};\bm{M})$. To demonstrate the impact of our proposed UNICORN framework, we discuss two instantiations as follows:
@@ 2.3 Instantiations of UNICORN
$I(\bm{Z};\bm{M})$ $=H(\bm{M})-H(\bm{M}|\bm{Z})\equiv-H(\bm{M}|\bm{Z})$
$=\mathbb{E}_{\bm{z}}\mathbb{E}_{M\sim p(M|\bm{z})}\left[\log p(M|\bm{z})\right]=-\mathbb{E}_{\bm{z}}\left[H(\bm{M}|\bm{Z}=\bm{z})\right].$ (7)
where $H(\cdot)$ is entropy. Since in practice, each $\bm{z}^{i}$ of sample $\bm{x}^{i}$ is collected within a specific task $M^{i}$, minimizing the parameterized entropy $H_{\bm{\theta}}(\bm{M}|\bm{Z}=\bm{z}^{i})$ is equivalent to finding an optimal function $p_{\bm{\theta}}(M|\bm{z})$ which correctly assigns the ground-truth label $M^{i}$ to $\bm{z}^{i}$, i.e., optimizing $p_{\bm{\theta}}(M|\bm{z})$ towards a delta function $\delta(M-M^{i})$ for continuous $\bm{M}$ or an indicator function $\mathbbm{1}(M=M^{i})$ for discrete $\bm{M}$. This implies that
@@
Suppose the task label $M$ is given during training and a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ for meta-training. Under this supervised scenario, by substituting Eqn 8 into 7, we have
@@
[…]
which is precisely the negative cross-entropy loss $H(\bm{M},P(\bm{M}|\bm{X}))$ for $n_{M}$-way classification with feature $\bm{z}$ and classifier $p_{\bm{\theta}}$. We therefore define the objective of supervised UNICORN as
$\mathcal{L}_{\textup{UNICORN-SUP}}$ $\coloneqq H(\bm{M},P(\bm{M}|\bm{X}))$
$=-\mathbb{E}_{\bm{x},\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\sum_{j=1}^{n_{M}}\mathbbm{1}(M^{i}=M)\log p_{\bm{\theta}}(M^{i}|\bm{z})\right].$ (10)
Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we give the following theorem:
Theorem 2.4 (Concentration bound for supervised UNICORN).
Denote by $\hat{I}(\bm{Z};\bm{M})$ the empirical estimate of $I(\bm{Z};\bm{M})$ by $n_{M}$ tasks, $\bar{I}(\bm{Z};\bm{M})$ the expectation, then with probability at least $1-\delta$,
$\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\leq\sqrt{\frac{\textup{Var}(H(\bm{Z}|\bm{M}))}{n_{M}\delta}}.$ (11)
@@
See Appendix B. ∎
Figure 3: Meta-learning procedure of UNICORN-SS. The supervised variant UNICORN-SUP simply replaces the decoder by a classifier $p_{\bm{\theta}}(M|\bm{z})$ and optimize a cross-entropy loss instead of $\mathcal{L}_{\textup{recon}}$ and $\mathcal{L}_{\textup{FOCAL}}$.
Theorem 2.4 bounds the finite-sample estimation error of the empirical risk $\hat{I}(\bm{Z};\bm{M})$ with $n_{M}$ task instances drawn from the real task distribution $p(\bm{M})$. The supervised UNICORN has the merit of directly estimating and optimizing the real objective $I(\bm{Z};\bm{M})$, which requires explicit knowledge of the task label $M^{i}$ and a substantial amount of task instances according to Theorem 2.4. To trade-off computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
Self-supervised UNICORN  In practice, offline RL datasets may often be collected with limited knowledge of the task specifications or labels. In this scenario, previous works implement self-supervised learning [31] to obtain effective representation $\bm{Z}$, such as the contrastive-based FOCAL/CORRO to optimize $I(\bm{Z};\bm{X})$/$I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ respectively; or generative approaches like VariBAD [32]/BOReL [13] to reconstruct the trajectories $\bm{X}$ by variational inference, which is equivalent to maximizing $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$. By Theorem 2.3, these methods optimize a relatively loose upper/lower bound of $I(\bm{Z};\bm{M})$, which can be improved by a convex combination of the two bounds:
@@
where $0\leq\alpha\leq 1$ is a hyperparameter. Implementing each term in Eqn 12 allows ample decision choices, such as the contrastive losses in Eqs. 3, 4 and 6 or autoregressive generation via Decision Transformer [33, 34] or RNN [32]. For demonstration, in this paper we employ a contrastive objective $\mathcal{L}_{\textup{FOCAL}}$ as in Eq. 3 for $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction. By the chain rule of mutual information [35]:
@@
$I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ $=\mathbb{E}_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\left[\log\frac{p(\bm{x_{t}}|\bm{z},\bm{x_{b}})}{p(\bm{x_{t}})}\right]$
$\equiv\mathbb{E}_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\left[\log p(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right]$
$\geq\mathbb{E}_{\bm{x_{t},x_{b}},\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}\left[\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right],$ (14)
which induces a generative objective $\mathcal{L}_{\textup{recon}}\coloneqq-I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$. As a result, the proposed unsupervised UNICORN objective can be rescaled as Eqn 15. The influence of the hyper-parameter $\frac{\alpha}{1-\alpha}$ is shown in Appendix C.4.
@@
$\mathcal{L}_{\textup{UNICORN-SS}}\coloneqq\mathcal{L}_{\textup{recon}}+\frac{\alpha}{1-\alpha}\mathcal{L}_{\textup{FOCAL}}.$ (15)
We summarize our learning procedure in Fig. 3 with pseudo-code in Algorithms 1 and 2. A holistic comparison of our proposed algorithms with related contextual meta-RL methods is shown in Table 1. The extra KL divergence can be interpreted as the result of a variational approximation to an information bottleneck [36, 37] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [38] is employed to tackle the bootstrapping error [39] for downstream offline RL implementation.
Table 1: Comparison between UNICORN instantiations and related existing contextual meta-RL methods. For clarity, "Representation Learning Objective" only lists the loss functions of $\bm{Z}$ that are independent of the downstream RL tasks. Note that $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\equiv I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ holds only for offline RL.
Method Setting Representation Learning Objective Implementation Context $\bm{X}$ UNICORN-SUP Offline $I(\bm{Z};\bm{M})$ Predictive Transition UNICORN-SS Offline $\alpha I(\bm{Z};\bm{X})+(1-\alpha)I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ Contrastive+Generative Transition FOCAL [11, 21] Offline $I(\bm{Z};\bm{X})$ Contrastive Transition CORRO [23] Offline $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ Contrastive Transition CSRO [22] Offline $(1-\lambda)I(\bm{Z};\bm{X})+\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ Contrastive Transition GENTLE [40] Offline $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ Generative Transition BOReL [13] Offline $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ Generative Trajectory VariBAD [32] Online $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ Generative Trajectory PEARL [7] Online $-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ N/A Transition ContraBAR [41] Offline&Online $I(\bm{Z};\bm{X}_{t}|\bm{A})$ Contrastive Trajectory